OpenStudy (anonymous):

solving matrix: 5x-y+2z=6 x+2y-z=-1 3x+2y-2z=1

OpenStudy (anonymous):

i know R1=5,-1,2,6 R2=1,2,-1,-1 R3=3,2,-2,1

OpenStudy (turingtest):

this is just painful to type out are you familiar with gaussian elimination?

OpenStudy (anonymous):

yes. i got z right. lol